A beam of light is traveling in air and strikes a material. The angles of incidence and refraction are 54.3° and 43.7°, respectively. Obtain the speed of light in the material.
Answer #1

Given,

theta-i = 54.3 deg

theta-r = 43.7 deg

We know from Snell's law

n1 sin(theta-i) = n2 sin(theta-2)

n1 = n(air) = 1

n2 = sin(theta-i)/sin(theta-r)

n2 = sin54.3/sin43.7 = 1.18

We know from the definition if index of refraction:

n = c/v

v = c/n

v = 3 x 10^8/1.18 = 2.54 x 10^8 m/s

Hence, v = 2.54 x 10^8 m/s